OpenStudy (anonymous):

8th grade math... Please help! The length of a rectangle is 8 in. more than its width. the perimeter of the rectangle is 24 in. what are the width and length of the rectangle.

OpenStudy (anonymous):

isnt it like 8x24?

OpenStudy (anonymous):

l = w + 8 p = 2l + 2w = 24 2(w+8) + 2w = 24 2w + 16 + 2w = 24 4w = 8 w = 2 l = 10

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

The length is 8in more than it's width, but they don't give the width. So call it w. Then what's the length (in terms of w)?

OpenStudy (anonymous):

hmm...